question 24 of chapter 10 rotation and rolling...///////////////////////////////

Grade:11

1 Answers

we will so
6 years ago
Moment of inertial of a square plate about its diagonal is ma2/12 (m = mass of the square plate)
a = edges of the square
Therefore torque produced = (ma2/12) × 
= {(120 × 10–3 × 52 × 10–4)/12 × 0.2
= 0.5 × 10–5 N-m.